Can you help me with solving: The total cost of producing x units is C=60x=300, whereas the revenue brought in of selling x units is R=100x-0.5x^2
How many units need to be sold to maximize reveune? what is the maximum revenue?

Answer by stanbon(75887)

Can you help me with solving: The total cost of producing x units is C=60x=300, whereas the revenue brought in of selling x units is R=100x-0.5x^2
How many units need to be sold to maximize reveune? what is the maximum revenue?
------
R(x) is a quadratic with a = -5 and b = 100
----
R(x) has a maximum value when x = -b/(2a) = -100/(2*-5) = 10
----
Maximum revenue = R(10) = 100*10 - 0.5(10^2) = 1000-50 = $950.00
===================
Cheers,
Stan H.
===================